OpenStudy (anonymous):

How many liters of 18.5 molar HCl stock solution will be needed to make 25.0 liters of 1.5 molar HCl solution? Show the work used to solve this problem.

OpenStudy (rogue):

Didn't Mert answer this in the math section? =P \[M_1 V_1 = M_2 V_2\]\[18.5 M \times V_1 = 1.5 M \times 25.0 L\]\[V_1 = \frac {37.5 mol HCl}{18.5M} = 2.03 L\]

OpenStudy (anonymous):

lol no that was a different one haha thank you(: